# Tilings from Tops of Overlapping Iterated Function Systems

### v2

@@ 1. Introduction
In [6] top addresses were introduced to define tilings in special cases. This paper elucidates the situation in [6] and answers in the affirmative the conjecture at the end of that paper. A distinct approach to overlapping tilings has been provided by Akiyama and co-workers [2].
@@ 2.4. Top blowups and tilings
In Theorem 1 we consider expressions of the form
$S(l,\mathbf{i},k):=f_{j_{1}}^{-1}\dots f_{j_{k}}^{-1}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}})$
where $k,l\geq 0$ and $\mathbf{i\in}\Omega_{l}$ with
$\Omega_{l}:=\left\{\mathbf{i}\in\Sigma:j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}\in\Sigma_{k+1}\text{ for all }k\geq l+1\right\}.$
Sets of the form $S(l,\mathbf{i},k)$ are supersets of pre-tiles that occur in the expression $\lim_{k\rightarrow\infty}\Pi(\mathbf{j}|k)$. The key observation underlying Theorem 1 is that, under a condition on $\mathbf{j}$, as $k$ increases the sequence of sets $S(l,\mathbf{i},k)$ decreases for finitely many steps after which it does not change.
@@
The tiling $\Pi(\overline{1})$ is well-defined and its support is the blowup ${\textstyle\bigcup\limits_{k}}f_{1}^{-k}(A).$ In this case the set of tiles is exactly ${\textstyle\bigcup\limits_{k}}f_{1}^{-k}(\{f_{\mathbf{i}|(k+1)}(A):$ $\mathbf{i\in}\Sigma\mathbf{\},}$ see also [6].
@@ 4.1. Leaf example of a two-dimensional top tiling
The overlapping attractor looks like a leaf, and is the union of two overlapping copies of itself, as illustrated in Figure 5. The point with address $\overline{\mathbf{1}}=111\dots$ is represented by the tip of the stem of the leaf. The stem is actually arranged in an infinite spiral, not visible in the picture.
@@ 6. Acknowledgement
We thank Louisa Barnsley for help with the artwork.

### v3

@@ 1. Introduction
In [6] a special case of tilings defined by top adddresses was introduced. This paper extensively generalizes and extends this earlier result.
A different approach to the construction of tilings using overlapping tiles has been provided by Akiyama and co-workers [2].
@@ 2.4. Top blowups and tilings
In Theorem 1 $\mathbf{j\in}\overleftarrow{\Sigma}$ is fixed throughout. We consider expressions of the form
$S(l,\mathbf{i},k):=f_{j_{1}}^{-1}f_{j_{2}}^{-1}\dots f_{j_{k}}^{-1}(A_{j_{k}j_{k-1}\dots j_{l+2}j_{l+1}i_{1}i_{2}\dots i_{l+1}})$
for fixed $l$ and successively larger values of $k,$ where $k\geq l+1\geq 1$ and $\mathbf{i\in}\Omega_{l}$ where
$\Omega_{l}:=\left\{\mathbf{i}\in\Sigma:j_{k}j_{k-1}\dots j_{l+2}j_{l+1}i_{1}i_{2}\dots i_{l+1}\in\Sigma_{k+1}\text{ for all }k\geq l+1\right\}.$
$\Omega_{l}$ is the set of addresses whose corresponding pre-tiles first appear at the $l$-th iteration and remain visible for the rest of the tiling construction.
The sets $S(l,\mathbf{i},k)$ are supersets of pre-tiles that occur in the expression $\lim_{k\rightarrow\infty}\Pi(\mathbf{j}|k)$. The key observation underlying Theorem 1 is that, under an extra condition on $\mathbf{j}$, as $k$ increases the sequence of sets $S(l,\mathbf{i},k)$ decreases for finitely many steps after which it does not change.
@@
The tiling $\Pi(\overline{1})$ is well-defined and its support is the blowup ${\textstyle\bigcup\limits_{k}}f_{1}^{-k}(A).$ In this case the entire tiling is provided by the collection of pretiles ${\textstyle\bigcup\limits_{k}}f_{1}^{-k}(\{A_{\mathbf{i}|(k+1)}:$ $\mathbf{i\in}\Sigma\mathbf{\},}$ because the sequence $\{f_{1}^{-k}(\{A_{\mathbf{i}|(k+1)}:$ $\mathbf{i\in}\Sigma\mathbf{\}\}}_{k=1}^{\infty}$ is nested increasing. The key point is that $\Omega_{l}=\Sigma_{l+1}$ and because of this, the blowup of every set in the top-partitioned attractor at depth $k$ is included in the final tiling. Examples of this kind were introduced in [6].
@@ 4.1. Leaf example of a two-dimensional top tiling
The attractor looks like a leaf, and is the union of two copies of itself, as illustrated in Figure 5. The proof that this attractor has nonempty interior is illustrated in Figure 8. The point with address $\overline{\mathbf{1}}=111\dots$ is represented by the tip of the stem of the leaf. The stem is actually arranged in an infinite spiral, not visible in the picture.
Figure 8. Illustration of the proof that the leaf attractor has nonempty interior. The white polygon is mapped by the IFS into two polygons whose interiors cover the interior of the original polygon.
@@ 6. Acknowledgement
We thank Louisa Barnsley for help with the artwork in general and for her substantial contribution to the proof that the leaf attractor has nonempty interior.